OpenStudy (anonymous):

Convert 3.275 to a mixed number. A) 3 11/400 B) 3 275/100 C) 3 100/275 D) 3 11/40

jimthompson5910 (jim_thompson5910):

3.275 = 3 + 0.275 3.275 = 3 + 275/1000 3.275 = 3 + 11/40 ... divide each part of the fraction by 25 3.275 = 3 11/40